OpenStudy (hippoharry):

An exterminator estimated there were 12,000 termites in a house. Each time the house was sprayed, the number of termites was reduced to one-fourth the previous number. How many termites were there after the house was sprayed x times? Write a function to represent this scenario. A. f(x)=12,000+(1/4)^x B. f(x)=1/4(12,000)^x C. f(x)= 12,000(1/4)^x D. f(x)=12,000 - (1/4)^x

OpenStudy (rolldextro):

The correct answer is C. You can try something like this. f(0)=12000 f(1)=3000 f(2)=750 f(3)=375/2 see the pattern? f(x)=(1/4)^x*12000